Grade 11Physical Chemistry

With what velocity should an ∝-particle travel towards the nucleus of a copper atom so as to arrive at a distance 10-13 metre from the nucleus of the copper atom?

Simran Bhatia
12 Years agoGrade 11

1 Answer

Aditi Chauhan
12 Years ago
As the ∝-particle travelling with velocity, ‘u’, stops at a distance 10-13 m, its K.E. becomes zero and gets converted into P.E.
∴ 1/2 mu2 = 1/4π ε0 * 2Ze2/r or u2 = Ze2/4πε0 m.r
Here, Z = 29 for Cu atom
=> u2 29 *(1.6 *10-19)2/3.14 *8.85 *10-12 *(4 *1.672 *10-27) *10-13
u = 6.3 * 106 m sec-1